[Milkymist-devel] M1rc3 production report
adam at sharism.cc
Wed Dec 28 23:19:27 PST 2011
* The whole M1rc3 run has been almost finished, a total yield ratio
88pcs / 90pcs. (97.8%) is came out. Remaining two boards are still under
While initial test of first board, we ran into darn of never booted up
about solutions from reset IC and diode since an unqualified and unverified
with strict reviews after rc2 run we made. Then thanks to Werner and
Sebastien's helps, this post-design verification has been implemented by
h/w patch called 'fix2b' .
* Classifications of failure:
- No VGA synchronized signals out :: ratio 27 / 90
caused by unqualified Schmitt-Trigger IC( u19 or u20 ), a whole rework of
replacing them on 90sets is done.
- Impedance/Current failure :: 9 / 90
1 / 9: unqualified Schmitt-Trigger IC( u7 ), a whole rework of replacing
u7 on 90 sets is done.
4 / 9: unqualified usb transceiver ( u16 or u17 ), rework of replacing
failed one is done.
1 / 9: by C30 or C36 ( not sure which one caused it since I removed them
at moment ) C0402 footprint, we'll need to review this footprint later.
1 / 9: by C154, C0805 footprint; the gap between two pads probably is
thinner, this is to be changed for sure.
1 / 9: by C104, C0805 footprint; this part is nearby area of rework of
reset circuit, and a solder splash on C104, is a victim of rework.
1 / 9: Impedance of TP2 to GND is 380 ohm that differs from normal 628
ohm. Quiescent current is still stays 0.54A in normal. Works well after
- Flash failure :: 11 / 90
Due to twice significant h/w patches( reworks ) called 'fix2' and
'fix2b', we have got total 11 pcs in fail from them. After used a version
of reflash script with 'lockflash', they are all fixed. see  in details.
This's especially thanks to Werner's found.
- VGA DDC :: 6 / 90
Records here indicates that (female to male) expendables connected to VGA
Monitor's bounded cable are prepared beforehand, so that frequently
replug-in and replug-out connections with m1 boards.
- Audio :: 2 / 90
by L1 is a 0402 package but footprint is 0603, needs to change 0402
- Ethernet ARP :: 1 / 90
Indicates that just plugged cable was contacted well though.
- Video I2C :: 1 / 90
The VIDEOIN_SDA doesn't still stay steadily at high level. Then used heat
air to blow AB17(VIDEO_SDA) ball and passed. It's probably caused by that
it's a flux stocked/mixed with volatile cleaner surrounding ball area while
several times of reworks nearby 'reset circuit' to cause it. A
infrastructure of clean way needs to be settle well.
- MIDI :: 12 / 90
10 / 12: unqualified Schmitt-Trigger IC( u7 ), a whole rework of
replacing u7 on 90 sets is done.
1 / 12: 0x7d, after reflashed successfully with './reflash_m1.sh --rc3'
version "2011-10-14" with locked bit, then TP31 works @ 3.3V
but still checked its level and worked well after overnight rendering test.
1 / 12: 0x46, failed by a 'foreign body' which mixed and composed of
unqualified process while printing soldering balls, see; after reflow by
bga rework station then became 3.3V level works well, see.
- USB :: 22 / 90
by unqualified usb transceiver, replaced the failure one and pass.
- Rendering :: 9 / 90
this recorded failure result( i.e. nor-corruption issue ) that in first
round of "fix2" rework was not effective. After rework of "fix2b" patch,
accept it and face it as pass.
- CRC :: 1 / 90
After applied fix2b circuit patch, solved.
* Improvements in topics next:
- A design verification being approved by enough relevant experiments.
- A specified and well-done temperature profile to be measured and best
that monitoring the reflow process in personally appearance. This is try to
do the smt works good at the beginning.
- Sourcing source picked must change to boom system or famous
- Reduce reworks as possible. If need, think it firstly before rework and
also make sure an infrastructure of clean way settled well on both factory
and my site.
* Known issues of M1rc3 main board run, see
- After then a nor corruption problem was discovered. A M1pre-rc4 proposed
reset circuit are came up.
- Video-In codec will be switched to 7181C in rc4.
- USB power switch can be added, Werner's test results of "USB current
limit vs. capacitative loads"
- replaced L19 with a short connection by soldering on all run3 boards to
solve the intermittent loss of video input.
- some topics like LEDs and DVI-I idea have been brought up.
* Some improvements to be prepared before rc4 about assembling and
With packing instructions, M1rc3 they've been assembled with acrylic
cases and packed with accessories.
- Incoming parts inspections :: cases, cleanliness of shield sheet, if
existed that big spacers without screw threads, insertion of retailer box.
- work out an efficient way to clean such "small bits and pieces" on
surface of rc3 top case after tearing off films.
If some details I didn't bring up with regardless here, they should have
been recorded in . I'll watch it out.
-------------- next part --------------
An HTML attachment was scrubbed...
More information about the Devel